if it takes my sister 2 hours to clean her room. and it takes me 1 hour and 30 minutes to clean mine. How long would it take if we worked together?
Answer by jorel1380(3719)

1/120+1/90=?
3/360+4/360=7/360
It would take 360/7 minutes working together to clean a room.